1. Calculate GDP in 2014 from the expenditure approach and advise which variable of GDP it is calculated under and advise what the final value of GDP is. Billions of dollars Government purchases of roads 10 Cars produced in Germany and 19 purchased in US Wheat purchased to manufacturer 4.5 bread Ford Motor Company putting in new 8.1 computers Unsold cars during 2014 (inventory) 5.3 Purchase of Stocks and Bonds Clothing made in China purchased 78 in US Groceries Cars produced in Ohio and sold to 2.8 Germany New construction of housing8.9 Government payments to Veterans urchase of 2008 Ford Explorer 5.2 urchase of refrigerators produce 7.5 in US 4 5.1 inventories from 2013 not used) 1.8 ayment to stockbrokers for ransfer of stocks and bonds urchase of 2013 Jeep (unsold
Answer #1

Gross Domestic product (GDP) is defined as the market value of all goods and services produced within the national boundary in a particular period of time. Here time generally referred to a quarter, or a year,as quarterly GDP or yearly GDP. Usually yearly GDP is calculated.

Economists have suggested three major approach to measure GDP of a country:

1). The product approach to measuring GDP. 2). The income approach to measuring GDP.   3). The expenditure approach to measuring GDP.

The expenditure approach measures GDP as total spending on final goods and services produced within a nation during a specific period of time. Four major components of spending are added to calculate GDP, also called equation for measuring GDP, i.e :

GDP = C+I+G+(X - M)

where, C = Private Consumption Expenditure, I= Private Investment Expenditure, G= Government Expenditure,

X = Exports from a particular country, M = Imports to a particular country,

Therefore, (X-M) is called Net Exports.

While calculating GDP we need to take care of double counting of value of the products.

Calculating GDP from the data given in above equation:

Most of the spending above which comes under four major components of GDP as C, I, G, X, M is included in measuring GDP.

GDP = Groceries+ Purchase of Refrigerator produces in US+ Payments to stock brokers + unsold cars during 2014 (inventories)+ New construction of housing + Government purchase of Roads + Ford motor company putting new computers + Cars produced in Ohio(A state in US) and sold to Germany - Cloth made in China and purchased in US - Cars produced in Germany and purchased in US

GDP = 9 + 7.5 + 4.4 + 5.3 +8.9 + 10 + 8.1 + 2.8 - 7.8 - 19 = 29.2

Wheat purchased for making bread is an intermediate good, not a final good so not included. Expenditure on financial products like bonds is not included as these also works as type of intermediate goods. Government payments to veterans is a unilateral transfer , is also not counted in GDP calculation. It is calculated using expenditure approach. The final value of GDP is 29.2 billion dollars.